OpenStudy (anonymous):

the least value of 4 cosec^2 A + 9 sin^2 A is: ? a)10 b)11 c)12 d)14

OpenStudy (kropot72):

I have got an answer by differentiating with respect to A and putting the result equal to zero. Then solving for A and substituting the value for A in the original expression.

OpenStudy (kropot72):

\[y=4\csc ^{2}A+9\sin ^{2}A\] \[\frac{dy}{dA}=(-8\csc ^{2}A \times \cot A)+18\sin A \cos A=0\]

OpenStudy (anonymous):

@dhiwagm use this hint \[4 \csc^2 A+9 \sin^2 A=\frac{4}{\sin^2 A}+9 \sin^2 A=\frac{4}{\sin^2 A}+9 \sin^2 A-12+12 \\ =(\frac{2}{\sin A}-3 \sin A)^2+12\] answer is 12